Wow. This should be easier...
AB * CD = DDD
What's D? There should be two answers, but I'm dumbfounded helping my daughter with her homework! :) Help

Found 2 solutions by haileytucki, nn-i:
Answer by haileytucki(390)

AB*CD=DDD
Since D is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
DDD=AB*CD
Combine all similar variables in the expression.
D^(3)=AB*CD
Multiply AB by CD to get ABCD.
D^(3)=ABCD
Since ABCD contains the variable to solve for, move it to the left-hand side of the equation by subtracting ABCD from both sides.
D^(3)-ABCD=0
Move all terms not containing D to the right-hand side of the equation.
-ABCD+D^(3)=0

Answer by nn-i(2)

There was a misunderstanding with the previous answer. For the question, A, B, C & D represent 4 different numbers and the multiplication of number AB with number CD results in a number DDD.

The key to solving this problem is to recognize that DDD has factors of D and 111.
In turn, 111 has factors 3 and 37.

Stop reading now as you may be able to solve with that clue.



Did you get it?

To continue (DDD) = (37 * 3 * D ) = (AB * CD)
Thus can infer that either AB or CD is 37 and the other number is D * 3.
Using guess and check to find D and we get 5 and 7 as possible value of D.
So, ensuring that the digit value of the second number is D ...
D=5:
37 * 3 * 5 = 555
37 * 15 = 555
A=3, B=7, C=1, D=5
and
D=7:
37 * 3 * 7 = 777
21 * 37 = 777
A=2, B=1, C=3, D=7